PREFACE

The 38th Symposium on Ring Theory and Representation Theory was held at Aichi Institute of Technology in Aichi Prefecture from September 2 to 4 in 2005. The symposium and the proceedings are financially supported by Toshiyuki Katsura (University of Tokyo) JSPS Grant-in-Aid for Scientific Research (A), No.15204001, Kenji Nishida (Shinsyu University) JSPS Grant-in-Aid for Scientific Research (C), No.17540021, and Masahisa Sato (Yamanashi University) JSPS Grant-in Aid for Scientific Research (C), No.16540019.

This volume consists of the articles presented at the symposium. We would like to thank all speakers and coauthors for their contributions.

We would also like to express our thanks to all the members of the organaizing committee for their helpful suggestions concerning the symposium. Finally we should like to express our gratitude to Professor Sumiyama and his students of Aichi Institute of Technology who contributed in the organization of the symposium.

Takayoshi Wakamatsu

Saitama, Japan March, 2006